OpenStudy (anonymous):

does -3/8 / (-1/6)= 9/4? explain.

OpenStudy (amistre64):

well the sign is right..... -/- = + now tell me, if we want to get rid of a denominator, what do we multiply by?

OpenStudy (anonymous):

hmm.. let me think it through

OpenStudy (anonymous):

either 2 or i thought you were suppose to, its on the top of my head but its not coming out

OpenStudy (amistre64):

-3/8 / (-1/6) we actually have 3 denominators to pick from, but lets use the bottom multiply the top and bottom by 6 -6(3/8) ------- -6(1/6)

OpenStudy (anonymous):

36(3/48) then simplify. right?

OpenStudy (amistre64):

-6(3/8) ------- -6(1/6) << 6/6 = 1 soo -6(3/8) ------- which is just 6(3/8) -1 which should be simple enough to reduce now

OpenStudy (anonymous):

oh okay that makes sense thanks :)

OpenStudy (amistre64):

instead of trying to remember some flip rule .... if you can clean out the fraction underneath it maths into something normal: \[\frac{a/b}{x/y}\] \[\frac{ya/b}{\cancel yx/\cancel y}\] \[\frac{ya/b}{x}\] \[\frac{y/x~a/b}{\cancel x/\cancel x}\] \[\frac{a}{b}~\frac{y}{x}\]

OpenStudy (amistre64):

the 'rule' is, flip the bottom and multiply

OpenStudy (anonymous):

hmm.... so the bottom number goes on the top? is what you are saying.

OpenStudy (amistre64):

im saying we can do the math and it works out as math is spose to work instead of trying to remember some rule that only applies to a specific situation

OpenStudy (anonymous):

OH

OpenStudy (amistre64):

the rule is .... flip the bottom and multiply \[rule:\frac{a/b}{x/y}\to~\frac ab \frac yx\]

OpenStudy (amistre64):

if you cant remember the rule .... so the math :)

OpenStudy (amistre64):

*do the math

OpenStudy (anonymous):

yeah i can remember the rule you just put better than the first lol

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

thank you for the help :)

OpenStudy (amistre64):

yw